OpenStudy (anonymous):

Solve 4y - y + 11 = 38. 9 -9 3.8 8.6

OpenStudy (anonymous):

3y+11=38

OpenStudy (anonymous):

Do u agree

OpenStudy (anonymous):

9

OpenStudy (anonymous):

@Viyang type a detailed explanation to make him/her understand

OpenStudy (anonymous):

3y+11=38

OpenStudy (anonymous):

Yeah...I don't get it..

OpenStudy (anonymous):

3y=38-11

OpenStudy (anonymous):

4y - y is 3y 38-11 is 27 27 / 3 is 9

OpenStudy (anonymous):

3y=27

OpenStudy (anonymous):

y=9

OpenStudy (anonymous):

@kbbbb @Viyang got it right

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

Do u get it?

OpenStudy (anonymous):

I understand better now

OpenStudy (anonymous):

Good